1) An organic compound X contains 40% carbon, 53.3% oxygen and 6.7% hydrogen. The vapour density of X is 30. The organic compound reacted in sodium hydrogen trioxocarbonate if and a colourless gas was given off which turn lime water milky. On boiling under reflux, the mixture of X and ethanol is a sweet smelling liquid Y was formed.

a) Determine the empirical and molecular formula of X.

b) What is the name of the colourless gas given off?

c) Given the formula of the functional group present in X.

d) Deduce the structure of the compound X.

e) Name the reaction which took place between X and ethanol and write an equation for the reaction between X and ethanol.


2) Calculate the number of moles of the material in the following:
2.25g of Na2co3
17.75g of Na2so4
100g of Naclo


3) Write the electronic configuration of:
Co^2+
W
Fe^3+
Zn^2+
Bi

And identify the group and period they belong to.
1
Expert's answer
2019-10-28T07:50:18-0400

1)   CxHyOz

      x : y : z = w(C)/Ar(C) : w(H)/Ar(H) : w(O)/Ar(O) = 40/12 : 6.7/1 : 53.3/16 = 3.33 : 6.7 : 3.33 ≈ 1 : 2 : 1.

a)   the empirical formula of X: CH2O.

Since the vapour density of X is 30, the molecular formula X is CH2O.

b)   By the condition of the problem, substance X is formaldehyde. But the described chemical properties of substance X say that X is formic acid. Then, the name of the colourless is carbon dioxide:

HCO2H + NaHCO3 = HCO2Na + H2O + CO2↑

     Ca(OH)2 + CO2 = CaCO3 + H2O

c) –COOH

d)


e) The Esterification Process:

HCO2H + C2H5OH = HCO2C2H5 + H2O



2)   n(Na2CO3) = m(Na2CO3)/M(Na2CO3) = 2.25g / 106g/mol = 0.021mol.

      n(Na2SO4) = m(Na2SO4)/M(Na2SO4) = 17.75g / 142g/mol = 0.125mol.

      n(NaClO) = m(NaClO)/M(NaClO) = 100g / 74.5g/mol = 1.342mol.


3)   Co^2+: 1s2 2s2 2p6 3s2 3p6 3d7

       W: 1s2 2s2 2p6 3s2 3p6 3d10 4s2 4p6 4d10 5s2 5p6 4f14 5d4 6s2

       Fe^3+: 1s2 2s2 2p6 3s2 3p6 3d5

       Zn^2+: 1s2 2s2 2p6 3s2 3p6 3d10

       Bi: 1s2 2s2 2p6 3s2 3p6 3d10 4s2 4p6 4d10 5s2 5p6 4f14 5d10 6s2 6p3